Circular No. 6814

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0054–720 J. C. Lochner, Goddard Space Flight Center (GSFC) and Universities

Space Research Association (USRA); F. E. Marshall, GSFC; L. A. Whitlock, GSFC and USRA; and N. Brandt, Pennsylvania State University, note: “RXTE detected a new pulsating object (period 169.30 s) on 1997 Dec. 17 in the direction of the Small Magellanic Cloud; the object has a 2–10-keV flux of 6.0 × 10^{−11} erg s^{−1} cm^{−2}. Scans made across the region on Dec. 20 revealed a single source located at α = 0^{h}54^{m}.6, δ = −72^{o}04′, (equinox 2000.0; estimated error circle 10′). By this time, the source had brightened to 8.3 × 10^{−11} erg s^{−1} cm^{−2}. Pulsations have persisted in observations taken through 1998 Jan. 12, with a monotonic decrease in the period to 168.40 s, and a decrease in the flux to 3.3 × 10^{−11} erg s^{−1} cm^{−2}. Pulsations from XTE J0053–724 (IAUC 6788, 6789) and AX J0051–722 (IAUC 6803) persist in a number of these observations. Simultaneous BeppoSAX and RXTE observations of the region are planned for Jan. 28.”

NGC 2060 G. Cusumano, M. C. Maccarone, T. Mineo, and B. Sacco, Istituto

di Fisica Cosmica con Applicazioni all’Informatica, CNR, Palermo, report: “Analysis of archival data acquired with the MECS instrument onboard the BeppoSAX X-ray Observatory confirms the RXTE detection of pulsations in NGC 2060 (IAUC 6810). BeppoSAX observed the LMC field on 1996 Oct. 25–26, and NGC 2060 was 16′ off-axis in the MECS field-of-view. In the energy band 2–10 keV, a pulsed signal was detected with a period of 0.016114775(4) s (in agreement with the extrapolated RXTE value); the folded lightcurve is has a single peak; the observed pulsed countrate, 0.0037 ± 0.0006 counts/s, corresponds to a flux of 5.9 × 10^{−13} erg cm^{−2} s^{−1}.”

COMET C/1995 O1 (HALE-BOPP) P. R. Holvorcem, Departamento de Matematica, Universidade Estad-

ual de Campinas, reports that the straight tail/antitail of C/1995 O1 reported on IAUC 6812 is visible faintly on an image obtained by co-adding three CCD exposures taken around Jan. 23.055 UT with a 0.20-m f/3.3 reflector. A relatively uniform distribution of faint comet material is bounded by the tail and antitail in p.a. ∼ 21^{o} and 197^{o}, respectively, covering its eastern side out to at least 15′ from the nucleus and overlapping most of the (much brighter) main dust tail. These images are available at

http://www.ime.unicamp.br/~holvorce.

1998 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green